mathslover (mathslover):

A charged oil drop with a mass of \(2 \times 10^{-4} \) kg is held suspended by a downward by an Electric field of 300 N/C . Find the charge on the drop.

mathslover (mathslover):

\(\color{blue}{\text{Originally Posted by}}\) @mathslover A charged oil drop with a mass of \(2 \times 10^{-4} \) kg is held suspended by a downward by an Electric field of 300 N/C . Find the charge on the drop. \(\color{blue}{\text{End of Quote}}\)

mathslover (mathslover):

@ganeshie8

OpenStudy (cwrw238):

sorry cant help you there

mathslover (mathslover):

No problem :)

OpenStudy (vincent-lyon.fr):

Weight is balanced by the electrostatic force. mg = |q|E => |q| = mg/E Now, if field is downward, since electrostatic force must be up, the q is negative. So final answer is : q = -mg/E

mathslover (mathslover):

Got it Sir. Thanks!